# Alejandro Benito-Santos, Adrián Ghajari Espinosa, Victor Fresno UNED research center in Natural Language Processing and Information Retrieval

### v1

@@ 6 Conclusions
RQ1: What is the performance of models in a dataset with minimal contamination? We have found that the behaviour of models is almost identical on our dataset and on MMLU-10: the model rankings provided by both datasets are almost equivalent (Pearson correlation is 0,98). In both datasets, performance is strongly correlated with model size. Across subjects, all models perform substantially worse in subjects requiring reasoning and calculus, compared to those which require mainly memorization. We plan to classify the questions according to the level of memorisation or reasoning they require, as in Yu et al. (2023) where they use Bloom’s taxonomy of educational objectives (Bloom, 1956) in order to better assess models on these dimensions. Overall, the fact that UNED-ACCESS has minimal contamination does not lead to any meaningful difference in the relative models’ performance. This may imply that for general knowledge subjects contamination is not crucial, or that there are indirect sources of contamination. In a preliminary investigation, a web search for solved exams in each of the subjects did not lead to any systematic repository with questions and answers, although we occasionally found PDF versions of the exams with the right answers highlighted with a pen (usually from academies that prepare the students for the exams).
@@
To mitigate contamination and increase difficulty, we plan to extend the dataset with undergraduate-level questions of increasing difficulty, and also to check the real level of contamination to which they may be exposed. We also plan to estimate the consistency of models’ answers to prompt variations, that is, how the models deal with changes in content (and get to measure a reasoning gap (Srivastava et al., 2024)), and in form (to measure the performance spread (Polo et al., 2024; Sclar et al., 2023; Mizrahi et al., 2024)), and apply this methodology to other benchmarks.
@@
Firstly, although the correct answers of the exam questions of UNED-ACCESS 2024 have never been made public officially, which makes us believe that the risk of contamination is limited, we have not measured the real level of contamination by, for example, making systematic searches in the internet in order to determine whether and in what form some correct answers may be publicly available. As stated in the paper, more research is needed to reliably estimate contamination. Secondly, the size of the dataset might be considered small as compared to the most popular public datasets. We have given priority to gather quality questions with private correct answers, rather than to build a large dataset. However, we are working towards producing a larger dataset with university level questions in a diversity of subjects. Third, our estimation of the language gap is just a first approximation, as there are confounding factors (translation, contamination) that are difficult to isolate from linguistic performance. More research is needed to find adequate methodologies to estimate the language gap. Also, our experimentation is directed at estimating a minimal gap (comparing English with other predominant online language such as Spanish); obviously, the gap will be larger with less resourced languages. Finally, we have not categorised the questions into levels of cognitive complexity, for example using Bloom’s taxonomy (Anderson and Krathwohl, 2001), in order to make an in-depth study of the models performance per cognitive level. This will be addressed in future work.
@@ Acknowledgments
This work has been funded by the European Union - NextGenerationEU through the ‘Recovery, Transformation and Resilience Plan’, by the Ministry of Economic Affairs and Digital Transformation and by UNED. However, the views and opinions expressed are solely those of the author(s) and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
@@ Appendix C Results
C.1 Results on UNED-ACCESS 2024
@@ C.2 Results on MMLU-10 by Subject
C.3 Checking Correlation of MMLU published Results and Results on MMLU-10
Table 7 provides the accuracy results on MMLU published in the official benchmark [^11] https://paperswithcode.com/sota/multi-task-language-understanding-on-mmlu (or, in their absence, in the respective model reports) and the results of applying our methodology on MMLU-10.
MMLU from publications MMLU-10 from zero-shot experiments
GPT-4o 88.7 83.40
Claude-3-Opus 86.8 (5-shot) 83.71
GPT-4-Turbo 86.5 81.61
Llama-3-70B-Instruct 82.0 Llama-3-70B (5-shot) 75.47
Gemma-2-27B-Instruct 75.2 Gemma-2 (5-shot) 76.61
Mixtral-8x7B-Instruct 70.6 Mixtral 8x7B (5-shot) 65.34
GPT-3.5-Turbo 70.0 63.74
Llama-3-8B-Instruct 68.4 Llama-3-8B (5-shot) 61.54
Mistral-7B-Instruct 58.4 (5-shot) 54.24
Gemma-7B-It 53.3 (5-shot) 42.50
Llama-2-7B-Chat 45.3 Llama-2-7B (5-shot) 35.08
Leniachat-Gemma-2B - 29.81
Pearson’s correlation between these two sets of results is 0.9863 with a p-value of 2.31e-08, which means that results are highly correlated. This high correlation validates our experimentation, as it implies that our choices of prompts and hyperparameters, and zero-shot setting, yield results consistent with those found in the literature.
@@ C.3 Checking Correlation of MMLU published Results and Results on MMLU-10
Table 6: Performance of models per subject on MMLU-10 (Cohen’s Kappa).

### v2

@@ 6 Conclusions
RQ1: What is the performance of models in a dataset with minimal contamination? We have found that the behaviour of models is almost identical on our dataset and on MMLU-10: the model rankings provided by both datasets are almost equivalent (Pearson correlation is 0,98). In both datasets, performance is strongly correlated with model size. Across subjects, all models perform substantially worse in subjects requiring reasoning and calculus, compared to those which require mainly memorization. We plan to classify the questions according to the level of memorisation or reasoning they require, as in Yu et al. (2023) where they use Bloom’s taxonomy of educational objectives (Bloom, 1956) in order to better assess models on these dimensions. Overall, the fact that UNED-ACCESS has minimal contamination does not lead to any meaningful difference in the relative models’ performance. This suggests that contamination affects similarly to all models, or that there are indirect sources of contamination in our dataset. In a preliminary investigation, however, a web search for solved exams in each of the subjects did not lead to any systematic repository with questions and answers, although we occasionally found PDF versions of the exams with the right answers highlighted with a pen (usually from academies that prepare the students for the exams).
@@
To further mitigate contamination and increase difficulty, we plan to extend the dataset with undergraduate-level questions of increasing difficulty, and also to check the real level of contamination to which they may be exposed. We also plan to estimate the consistency of models’ answers to prompt variations, that is, how the models deal with changes in content (and get to measure a reasoning gap (Srivastava et al., 2024)), and in form (to measure the performance spread (Polo et al., 2024; Sclar et al., 2023; Mizrahi et al., 2024)), and apply this methodology to other benchmarks.
Finally, although the questions and answers will not be distributed to keep contamination minimal, new relevant open models will continue to be evaluated in a UNED-ACCESS leaderboard at https://leaderboard.odesia.uned.es. Developers with an open model that wish to be included in the evaluation may send a request to odesia-comunicacion@lsi.uned.es.
@@
Firstly, although the correct answers of the exam questions of UNED-ACCESS 2024 have never been made public officially, which makes us believe that the risk of contamination is limited, we have not measured the real level of contamination by, for example, making systematic searches in the internet in order to determine whether and in what form some correct answers may be publicly available. As stated in the paper, more research is needed to reliably estimate contamination. Secondly, the size of the dataset might be considered small as compared to the most popular public datasets. We have given priority to gather quality questions with private correct answers, rather than to build a large dataset. However, we are working towards producing a larger dataset with university level questions in a diversity of subjects. Third, our estimation of the language gap is just a first approximation, as there are confounding factors (translation, contamination) that are difficult to isolate from linguistic performance. More research is needed to find adequate methodologies to estimate the language gap. Also, our experimentation is directed at estimating a minimal gap (comparing English with other predominant online language such as Spanish); obviously, the gap will be larger with less resourced languages. Finally, we have not categorized the questions into levels of cognitive complexity, for example using Bloom’s taxonomy (Anderson and Krathwohl, 2001), in order to make an in-depth study of the models performance per cognitive level. This will be addressed in future work.
@@ Acknowledgments
This work has been funded by the European Union - NextGenerationEU through the ‘Recovery, Transformation and Resilience Plan’, by the Ministerio para la transformación digital y de la función pública and by UNED via cooperation agreement C039-21OT. However, the views and opinions expressed are solely those of the author(s) and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
@@ Appendix C Results
C.1 Results on UNED-ACCESS 2024 by Subject
@@ C.2 Results on MMLU-10 by Subject
Table 7: Cohen’s Kappa performance of models per subject on MMLU-10.